# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v2

@@ 2.2.2 The Restriction and Lifting Operators
The above definitions of the restriction and lifting operators guarantee that the reconstructed, by the lifting operator, normalized density profiles preserve the total density within the domain $\Omega$, as per Proposition 1. We highlight here that while the operators are constructed based on the dataset $X$, their employment applies to any seen or unseen observation $x(t)\in\mathbb{R}^{n_{c}}$ obtained from Eq. (19) and $\boldsymbol{y}(t)\in\mathbb{R}^{d}$.
@@ 2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
With the restriction operator, which maps density fields to latent coordinates, constructed, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data. We note that while the restriction and lifting operators $\mathcal{R}$ and $\mathcal{L}$ are, in principle, defined over continuous time inputs, we apply them at a discrete set of uniformly spaced time instances $t_{k}=t_{0}+k\delta t$ for $k=1,2,\ldots,K$, consistent with the temporal sampling of our data; see Eq. (8). To consider lags in the latent coordinate embeddings $\boldsymbol{y}(t_{k})$, we employ both linear and nonlinear autoregressive models, namely MVARs and LSTMs.
Let an MVAR model of order $w$ (i.e., with $w$ time lags) be expressed as:
$\boldsymbol{y}_{c}(t_{k})=\boldsymbol{A}_{0}+\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j})+\boldsymbol{\varepsilon}(t_{k}),$ (22)
where $\boldsymbol{y}_{c}(t_{k})$ is the latent variable vector, evaluated at discrete times $t_{k}$ for each of the $c=1,\ldots,C$ cases of different initial condition considered, $\boldsymbol{A}_{0}\in\mathbb{R}^{d}$ is the intercept vector and $\boldsymbol{A}_{j}\in\mathbb{R}^{d\times d}$ are the regression coefficient matrices for each lag $j$. The predictor vectors $\boldsymbol{y}_{c}(t_{k-j})\in\mathbb{R}^{d}$ include the $w$ past lag latent variables evaluated at discrete times $t_{k-1},\ldots,t_{k-w}$, while the error term $\boldsymbol{\varepsilon}(t_{k})\in\mathbb{R}^{d}$ is assumed to follow a multivariate white noise process:
@@
Given sequential (time series) data for $\{\boldsymbol{y}_{c}(t_{k})\}$, the unknown parameters of the MVAR in Eq. (22), i.e., the intercept vector $\boldsymbol{A}_{0}$ and the coefficient matrices $\boldsymbol{A}_{j}$, can be computed by the solution of the following least-squares problem:
$\operatorname*{argmin}_{\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}}\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right),$ (24)
@@
$\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\sum_{c=1}^{C}\sum_{k=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert_{2}^{2},$ (25)
@@
$\hat{\boldsymbol{y}}_{c}(t_{k})=\boldsymbol{A}_{0}+\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j}),$ (26)
for $c=1,\ldots,C$ and $k=1,\ldots,K$. The optimization problem in Eq. (24) can be solved using methods such as regularized least squares; in particular, we used ridge regularization [86]. We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations. If the latent coordinate time series is non-stationary, pre-processing steps like differencing can be applied to achieve stationarity before solving the optimization problem. We highlight that MVAR model training further requires the selection of an optimal lag window size $w$, which determines the length of the time series history used to model the temporal dependencies. Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in Appendix D.
As a non-linear autoregressive model alternative to linear MVARs, we further consider LSTM networks, a type of Recurrent Neural Networks (RNNs) characterized by gating mechanisms, which allow the network to retain and propagate information over long sequences selectively, but also effectively handle vanishing gradients; a common challenge in traditional RNNs [87, 88, 89, 90]. Let an LSTM model with past $w$ latent states $\boldsymbol{y}_{c}(t_{k-j})$ for $j=1,\ldots,w$ be expressed as:
@@
where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$. The gating mechanism of the LSTM is included in Eq. (27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for conciseness of the presentation; a vanilla LSTM including 4 gates, and contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
@@
where the LSTM loss function is the mean squared error (MSE) between the true value of the latent variable vector $\boldsymbol{y}_{c}(t_{k})$ and the predicted one $\hat{\boldsymbol{y}}_{c}(t_{k})$, as provided by Eqs. (27, 28), that is:
$\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert^{2}.$ (30)
To solve the optimization problem in Eq. (30), one typically uses gradient-based methods; in particular, we employed the Adam optimizer [91], which is a stochastic variant of gradient descent, known for its adaptive learning-rate capabilities.
@@
Finally, the predicted latent coordinates $\hat{\boldsymbol{y}}_{c}(t_{k})$ obtained from the trained MVARs or LSTMs models are subsequently mapped back to the high-dimensional density space using the lifting operator defined in Eq. (21).
3 Case study and methodology implementation
3.1 Configuration
The proposed methodology is employed on a crowd dynamics configuration where pedestrians are moving from left to right in a corridor of dimensions $48\,\text{m}\times 12\,\text{m}$ ($x$: length, $y$: width) in the presence of an obstacle, as shown in Fig. 2.
Figure 2: The case study: The SFM simulating pedestrians moving in a corridor from the left to the right with an obstacle.
The square obstacle of area $12.96\,\text{m}^{2}$ is centered at $(25.8\,\text{m},1.8\,\text{m})$, with the pedestrians navigating around it towards reaching the corridor’s right end. Hence, the computational domain is $\Omega=\left([0,48]\times[0,12]\right)\backslash\left([24,27.6]\times[0,3.6]\right)$.
For modelling the pedestrians’ movement, we use the SFM [1] as a microscopic agent-based simulator. SFM falls within the general form of Eqs. (1, 2) and the idea behind it is that the motion of individuals is governed by “virtual forces”. These forces include acceleration toward the desired destination, as well as repulsive forces “exerted” by obstacles, borders, and nearby individuals to maintain a safe distance. Additionally, one may also consider an attractive force due to mimetic behavior, which can be observed in scenarios such as emergency evacuations or collective movement toward objects of common interest [1]. The derivation of SFM is based on the Langevin-type equation that contains a frictional term, a deterministic force term, and a random force term. In the SFM, these terms translate into a goal-directed force (playing the role of friction), social and repulsive (from obstacles) deterministic forces, and a random term, reading:
@@ 3.1 Configuration
where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time. Here, we avoid the inclusion of random forces, since we are interested in the macroscopic behavior that often emerges from the deterministic pedestrian interactions [1]. Details on the SFM forces included in Eq. (31) and the related parameters that were considered in this work are provided in Appendix A.
We simulate a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see in the Appendix B). Pedestrians are directed to move from left to right by setting their initial target at coordinates $(x,y)=(25\,\text{m},\,9.6\,\text{m})$, corresponding to the midpoint of the corridor in the horizontal direction and 80% of its vertical extent. Upon reaching this intermediate waypoint, their final destination is updated to $(x,y)=(48\,\text{m},\,4.2\,\text{m})$, which lies at the right end of the corridor and at 35% of its height. To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents at the left boundary at $x=0\,\text{m}$, when agents reach the right boundary at $x=48\,\text{m}$.
Pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator. Such a configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.

### v3

@@ 2.2.2 The Restriction and Lifting Operators
In the case of two interacting populations with normalized densities $x^{(l)}(t)$ for $l=1,2$, we consider the augmented bases $\tilde{U}^{(l)}$ in Proposition 2 and adjust the restriction operator $\mathcal{R}$ to:
$\boldsymbol{y}(t)=[\boldsymbol{y}^{(1)}(t),\boldsymbol{y}^{(2)}(t)]^{\top}\in\mathbb{R}^{d_{1}+d_{2}+2m+2},\quad\boldsymbol{y}^{(l)}(t)=\mathcal{R}^{(l)}(x^{(l)}(t))=(\tilde{U}^{(l)})^{\top}\left(x^{(l)}(t)-\frac{1}{n_{t}}X^{(l)}\mathbf{1}_{n_{t}}\right)\in\mathbb{R}^{d_{l}+m+1},$ (36)
and the lifting operator $\mathcal{L}$ to:
$x(t)=[x^{(1)}(t),x^{(2)}(t)]^{\top}\in\mathbb{R}^{2n_{c}},\quad x^{(l)}(t)=\mathcal{L}(\boldsymbol{y}^{(l)}(t))=\tilde{U}^{(l)}\boldsymbol{y}^{(l)}(t)+\frac{1}{n_{t}}X^{(l)}\mathbf{1}_{n_{t}}\in\mathbb{R}^{n_{c}}.$ (37)
In both cases, the above definitions of the restriction and lifting operators guarantee that the reconstructed normalized density profiles preserve the total per-population density within the domain $\Omega$, as per Propositions 1 and 2. We highlight here that while the operators are constructed based on the dataset $X$ (or $X^{(l)}$ for $l=1,2$), their employment applies to any seen or unseen observation $x(t)$ obtained from Eq. (33) and $\boldsymbol{y}(t)$. Hereafter, for conciseness of the presentation, we will carry the single-population dimensions of $x(t)$ and $\boldsymbol{y}(t)$, unless otherwise specified.
Remark 2.1.
Mass preservation and element-wise positivity in the reconstructed field are distinct properties. Standard POD/SVD truncation does not guarantee non-negativity in reconstruction: negative pointwise values may appear even when the original snapshots are positive. Since the reconstructed field is obtained by adding back the original mean, small negative entries—when they occur—are expected to appear in the low-density tails of the distribution. In such cases, we post-process the corresponding snapshots only for observation purposes, by setting these values to zero, and renormalizing the reconstruction by the total mass of the nonnegative part. This ensures positivity while preserving total mass, with normalization factors remaining near 1.
Remark 2.2.
The reconstruction step is only observational in our pipeline, as density fields are reconstructed via $\mathcal{L}$ in Eq. (35) or (37), while the ROM evolves entirely in the latent space (see Eq. (5)). Since learning is performed with the original snapshots (which are nonnegative by construction), any post-process non-negativity enforcement does not influence the latent representation learning. Constrained reconstruction approaches, such as nearest-neighbor based methods commonly used in Diffusion Maps latent spaces [93, 26], can guarantee non-negativity by construction but require higher computational cost; integrating such techniques into our pipeline is straightforward, although a detailed comparative study is out of the scope in the current work.
@@ 2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
With the restriction operator, which maps density fields to latent coordinates, constructed, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data. We note that while the restriction and lifting operators $\mathcal{R}$ and $\mathcal{L}$ are, in principle, defined over continuous time inputs, we apply them at a discrete set of uniformly spaced time instances $t_{k}=t_{0}+k\delta t$ for $k=1,2,\ldots,K$, consistent with the temporal sampling of our data; see Eq. (8). To capture temporal dependencies via delayed embeddings (in line with Takens’/Whitney’s embedding theorems [106, 99]), we employ autoregressive ROMs to learn the dynamics of the latent embeddings $\boldsymbol{y}(t_{k})$. The simplest such model is MVAR, which provides an efficient, linear baseline with globally optimal, least squares solutions. For a nonlinear comparison, we employ LSTM networks [64], whose gating mechanisms effectively model complex temporal patterns; though numerous alternatives exist (Transformers [60], Neural Ordinary Differential Equations (ODEs) [98], etc.), LSTMs represent state-of-the-art in nonlinear sequential modeling. A single ROM is trained in both single- and two-population cases; in the latter, this supports coupled latent dynamics that capture temporal interactions between groups. For compactness in what follows, we denote the dimension of the concatenated two-population latent coordinates in Eq. (36) by $d=d_{1}+d_{2}+2(m+1)$, to match the single-population dimensions.
Let an MVAR model of order $w$ (i.e., with $w$ time lags) without intercept be expressed as:
$\boldsymbol{y}_{c}(t_{k})=\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j})+\boldsymbol{\varepsilon}(t_{k}),$ (38)
where $\boldsymbol{y}_{c}(t_{k})$ is the latent variable vector, evaluated at discrete times $t_{k}$ for each of the $c=1,\ldots,C$ cases of different initial condition considered, and $\boldsymbol{A}_{j}\in\mathbb{R}^{d\times d}$ are the regression coefficient matrices for each lag $j$. The predictor vectors $\boldsymbol{y}_{c}(t_{k-j})\in\mathbb{R}^{d}$ include the $w$ past lag latent variables evaluated at discrete times $t_{k-1},\ldots,t_{k-w}$. Since the latent variables $\boldsymbol{y}_{c}(t_{k})$ are zero-mean (a consequence of the mean-centered POD projection) the intercept term is excluded from Eq. (38). The error term $\boldsymbol{\varepsilon}(t_{k})\in\mathbb{R}^{d}$ is assumed to follow a multivariate white noise process:
@@
Given sequential (time series) data for $\{\boldsymbol{y}_{c}(t_{k})\}$, the unknown parameters of the MVAR in Eq. (38), i.e., and the coefficient matrices $\boldsymbol{A}_{j}$, can be computed by the solution of the following least-squares problem:
$\operatorname*{argmin}_{\{\boldsymbol{A}_{j}\}_{j=1}^{w}}\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right),$ (40)
@@
$\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\sum_{c=1}^{C}\sum_{k=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert_{2}^{2},$ (41)
@@
$\hat{\boldsymbol{y}}_{c}(t_{k})=\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j}),$ (42)
for $c=1,\ldots,C$ and $k=1,\ldots,K$. The optimization problem in Eq. (40) can be solved using regularized least squares methods [65]. We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations. If the latent coordinate time series is non-stationary, pre-processing steps like differencing can be applied to achieve stationarity before solving the optimization problem. We highlight that MVAR model training further requires the selection of an optimal lag window size $w$, which determines the length of the time series history used to model the temporal dependencies. Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in D.
As a non-linear autoregressive model alternative to linear MVARs, we further consider LSTM networks [64], a type of Recurrent Neural Networks (RNNs) characterized by gating mechanisms, which allow the network to retain and propagate information over long sequences selectively, but also effectively handle vanishing gradients; a common challenge in traditional RNNs [56, 25, 113]. Let an LSTM model with past $w$ latent states $\boldsymbol{y}_{c}(t_{k-j})$ for $j=1,\ldots,w$ be expressed as:
@@
where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$. The gating mechanism of the LSTM is included in Eq. (43) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for conciseness of the presentation; a detailed presentation of the gating mechanism in $F_{\mathrm{LSTM}}(\cdot)$ and the related parameters $\boldsymbol{\theta}$ can be found in [64, 56]. Here, we use a vanilla LSTM with 4 gates, containing the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
@@
where the LSTM loss function is the mean squared error (MSE) between the true value of the latent variable vector $\boldsymbol{y}_{c}(t_{k})$ and the predicted one $\hat{\boldsymbol{y}}_{c}(t_{k})$, as provided by Eqs. (43, 44), that is:
$\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert^{2}_{2}.$ (46)
To solve the optimization problem in Eq. (46), one typically uses gradient-based methods. Here, we employ the Adam optimizer [74], which is a stochastic variant of gradient descent, known for its adaptive learning-rate capabilities.
@@
Finally, the predicted latent coordinates $\hat{\boldsymbol{y}}_{c}(t_{k})$ obtained from the trained MVARs or LSTMs models are subsequently mapped back to the high-dimensional density space using the lifting operator defined in Eqs. (35) or (37) for the single or two-population case, respectively.
3 Case studies and methodology implementation
We considered two representative benchmark problems in crowd dynamics: a unidirectional crowd flow moving through a corridor containing an obstacle, and a counterflow of two interacting populations in the same domain. The configurations of both case studies are described below, along with the implementation of the proposed methodology.
3.1 Configurations
The two crowd dynamics configurations simulate pedestrians moving in a corridor of dimensions $48\,\text{m}\times 12\,\text{m}$ ($x$: length, $y$: width) in the presence of an obstacle, as shown in Fig. 2.
(a) Unidirectional flow
(b) Counterflow
Figure 2: Crowd dynamics configurations of SFM simulating pedestrians moving in a corridor past an obstacle. In the unidirectional flow configuration in Fig. 2(a), pedestrians move from the left to the right of the corridor. In the counterflow configuration in Fig. 2(b), two pedestrian populations move in opposite directions; group 1 (blue) moves from the left to the right, while group 2 (red) follows the opposite direction.
The square obstacle of area $12.96\,\text{m}^{2}$ is centered at $(25.8\,\text{m},1.8\,\text{m})$, with the pedestrians navigating around it towards reaching their destination (right or left corridor’s boundary). Hence, the computational domain is $\Omega\backslash\Gamma$, with the corridor occupying $\Omega=[0,48]\times[0,12]$ and the obstacle $\Gamma=[24,27.6]\times[0,3.6]$.
For modelling the pedestrians’ movement, we used the SFM [62, 61] as a microscopic agent-based simulator. SFM falls within the general form of Eqs. (1, 2) and the idea behind it is that the motion of individuals is governed by “virtual forces”. These forces include acceleration toward the desired destination, as well as repulsive forces “exerted” by obstacles, borders, and nearby individuals to maintain a safe distance. Additionally, one may also consider an attractive force due to mimetic behavior, which can be observed in scenarios such as emergency evacuations or collective movement toward objects of common interest [62]. The derivation of SFM is based on the Langevin-type equation that contains a frictional term, a deterministic force term, and a random force term. In the SFM, these terms translate into a goal-directed force (playing the role of friction), social and repulsive (from obstacles) deterministic forces, and a random term, reading:
@@ 3.1 Configurations
where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time. Here, we avoid the inclusion of random forces, since we are interested in the macroscopic behavior that often emerges from the deterministic pedestrian interactions [62]. Details on the SFM forces included in Eq. (47) and the related parameters that were considered in this work are provided in A.
Remark 3.1.
Here, we employed the Social Force model purely as a convenient baseline to demonstrate and validate the proposed methodology. Naturally, the same workflow is not tied to this specific choice: it can be coupled just as well with more sophisticated agent-based approaches—e.g., models that include explicit route planning, richer interaction rules, heterogeneous behaviors, or learned decision-making—depending on the fidelity required by the application.
In the unidirectional flow configuration (see Fig. 2(a), we simulated a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see C). Pedestrians are directed to move from the left to the right of the corridor. Since the SFM governs pedestrians’ motion through target-directed forces without intrinsic path planning, we employ a sequential two-target scheme to induce realistic obstacle-avoidance behavior. Pedestrians are initially guided towards an intermediate target zone below the obstacle at coordinates $x=25\,\text{m}$ and $y\in[6,7.8]\,\text{m}$. Upon reaching this zone, their final destination is updated to $x=48\,\text{m}$ with $y$-coordinates remaining at the values each agent had upon reaching the intermediate zone. To maintain a constant population density, we implemented periodic boundary conditions by re-spawning agents at the left boundary at $x=0\,\text{m}$, when agents reach the right boundary at $x=48\,\text{m}$.
In the counterflow configuration (see Fig. 2(b), we simulated a crowd of $N=100$ pedestrians divided equally into two groups with opposing flow directions: group 1 moves from left to right, and group 2 from right to left. We used the SFM in Eq. (47) with the same parameters as in the unidirectional case, and initial spatial distributions that are symmetric for each group along $x=24\,\text{m}$; group 1 follows the same initial distribution as in the unidirectional case. Employing a two-target scheme to induce realistic obstacle-avoidance behavior, we prescribed distinct preferred zones for each group. Group 1 is initially guided to the zone $x=25\,\text{m}$ and $y\in[5.4,7.8]\,\text{m}$, while group 2 to $x=25\,\text{m}$ and $y\in[9.6,11.4]\,\text{m}$. Upon reaching these zones, their final destination is updated to $x=48\,\mathrm{m}$ for group 1 and $x=0\,\mathrm{m}$ for group 2, while the $y$-coordinates remain fixed at the values each agent had upon reaching the intermediate zone. This separation promotes smooth obstacle avoidance and organized counterflow without introducing explicit route planning. As in the unidirectional case, periodic boundary conditions are applied at both ends of the corridor by re-spawning agents at the left/right boundary upon exiting at the right/left for groups 1/2, respectively.
In both configurations pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator. Such a configuration ensures a controlled pedestrian flow through the simulated environment and a realistic obstacle-avoidance behavior.